OpenStudy (anonymous):

Build the fraction to an equivalent fraction with the specified denominator. State the numerator 6/13 ?/65.

OpenStudy (king):

30/65

OpenStudy (king):

numerator is 30

OpenStudy (anonymous):

6*5/13*5 = 30/65

hero (hero):

6/13 = 30/65